# On anomalous dissipation induced by transport noise

### v2

@@ A.2. Stochastic Meyers’ estimates for the turbulent Stokes system
where we norm the product space ${\mathbb{R}}^{d}\times{\mathbb{R}}^{d\times d}\times\ell^{2}$ with the norm $\|(\cdot,\cdot,\cdot)\|_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d\times d}\times\ell^{2}}^{2}=\|\cdot\|_{{\mathbb{R}}^{d}}^{2}+\|\cdot\|_{{\mathbb{R}}^{d\times d}}^{2}+\|\cdot\|_{\ell^{2}}^{2}$. The latter choice comes from the fact that $\|\mathcal{S}_{2,2}\|_{\mathscr{L}}\leq 1$ as we have proved that (A.22) holds with $K_{2}=1$. The well-definiteness of $\mathcal{S}_{q,p}$ follows from the above-noticed existence and uniqueness of strong solutions $v\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ of (A.12) with data $g\in L^{p}_{\mathscr{P}}({\mathbb{R}}_{+}\times\Omega;\mathbb{L}^{q}(\mathbb{T}^{d};\ell^{2}))$ due to Step 1. Thus the claim of Substep 2a follows by complex interpolation [51, Theorem 2.2.6] and the elementary inequality $a^{\alpha}+b^{\alpha}\leq(a+b)^{\alpha}\leq 2^{\alpha-1}(a^{\alpha}+b^{\alpha})$ for $\alpha>1$.
Substep 2b: Proof of Theorem A.2. In light of Substep 2a, the estimate (A.21) with constant independent of $\lambda\in[0,1]$ and $f\equiv 0$, now readily follows from $\kappa>\kappa_{0}$ and the boundedness of $(\xi_{n})_{n\geq 1}$, see (A.13). ∎
@@ Acknowledgements
The author thanks Sebastian Bechtel, Nicolas Clozeau and Max Sauerbrey for useful suggestions and discussions. The author is grateful to Franco Flandoli for his suggestions on the heuristic derivation of Subsection 1.2. Finally, the author thanks Umberto Pappalettera for fruitful discussions and several comments which improved considerably the presentation of the results.

### v3

@@ A.2. Stochastic Meyers’ estimates for the turbulent Stokes system
where $v$ is the strong solution to (A.26) with $\lambda=0$ and $f\equiv 0$, which exists due to the comment below (A.28). We emphasize that the norm on product space ${\mathbb{R}}^{d}\times{\mathbb{R}}^{d\times d}\times\ell^{2}$ is the Euclidean one: $\|(\cdot,\cdot,\cdot)\|_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d\times d}\times\ell^{2}}^{2}=\|\cdot\|_{{\mathbb{R}}^{d}}^{2}+\|\cdot\|_{{\mathbb{R}}^{d\times d}}^{2}+\|\cdot\|_{\ell^{2}}^{2}$. The latter choice comes from the fact that $\|\mathcal{S}_{2,2}\|_{\mathscr{L}}\leq 1$ as we have proved that (A.28) holds with $K_{2}=1$. The well-definiteness of $\mathcal{S}_{q,p}$ follows from the above-noticed existence and uniqueness of strong solutions $v\in L^{p}({\mathbb{R}}_{+}\times\Omega;H^{1,q})$ of (A.20) with data $g\in L^{p}_{\mathscr{P}}({\mathbb{R}}_{+}\times\Omega;\mathbb{L}^{q}(\mathbb{T}^{d};\ell^{2}))$ due to Step 1. Thus, the claim of Substep 2a follows by complex interpolation [57, Theorem 2.2.6].
Substep 2b: Proof of Theorem A.3. In this substep, we prove the a priori estimate (A.27) for strong solutions to (A.26). Now, let $v_{\lambda}\in L^{p}((0,\tau)\times\Omega;\mathbb{H}^{1,q})$ be a strong solution to (A.26) where $\tau$ is a given stopping time. From the estimate of Substep 2a and the elementary inequalities $(a^{r}+b^{r})\leq(a+b)^{r}\leq 2^{r-1}(a^{r}+b^{r})$ that is valid for all $r\geq 1$ and $a,b\geq 0$, we obtain the existence of a constant $N_{p}$ such that $\lim_{p\downarrow 2}N_{p}=1$ such that
$\|v_{\lambda}\|_{L^{p}((0,\tau)\times\Omega;H^{1,q})}^{2}+\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}])\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2}$
$\leq N_{p}\Big\|\Big[(\kappa-\kappa_{0})(|v_{\lambda}|^{2}+|\nabla v_{\lambda}|^{2})+\|(\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}]\big)_{n\geq 1}\big\|^{2}_{\ell^{2}}\Big]^{1/2}\Big\|_{L^{p}((0,\tau)\times\Omega;L^{q})}^{2}$
$\leq N_{p}K_{p}\|(g_{n}+\lambda\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}])_{n\geq 1}\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2}$
$\leq N_{p}K_{p}(1+\delta^{-1})\|g\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2}$
$\quad+N_{p}K_{p}(1+\delta)\|\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}])_{n\geq 1}\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2},$
where $\delta>0$ is a positive constant, which will be fixed later. To conclude, it remains to show that the last term can be absorbed on the LHS of the corresponding estimate. To this end, from (A.17), it follows that, for all $p\in[2,p_{1}]$ and $q\in[2,p]$ (here $p_{1}$ is as in Substep 2a),
$N_{p}K_{p}(1+\delta)\|\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}])_{n\geq 1}\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2}$
$\leq\|\mathbb{P}[(\xi_{n}\cdot\nabla)v_{\lambda}])_{n\geq 1}\|_{L^{p}((0,\tau)\times\Omega;L^{q}(\ell^{2}))}^{2}+(1-N_{p}K_{p}(1+\delta))C_{p}M_{0}\|v_{\lambda}\|_{L^{p}((0,\tau)\times\Omega;H^{1,q})},$
where $C_{p}=\sup_{2\leq q\leq p}\|\mathbb{P}\|_{\mathscr{L}(L^{q})}$. Note that, by interpolation and the fact that $\|\mathbb{P}\|_{\mathscr{L}(L^{2})}=1$ (as $\mathbb{P}$ is an orthogonal projection on $L^{2}$), it follows that $C_{p}\to 1$ as $p\downarrow 2$. Hence, the claimed a priori estimate (A.27) for $v_{\lambda}$ follows by collecting the previous inequalities, and choosing $p\in[2,p_{1}]$ and $\delta>0$ such that $(1-N_{p}K_{p}(1+\delta))C_{p}M_{0}<\kappa-\kappa_{0}$. Note that the latter choice is possible as $\lim_{p\downarrow 2}N_{p}=\limsup_{p\downarrow 2}K_{p}=1$, where the latter follows from Substep 2a. ∎
@@ Acknowledgements
The author thanks Sebastian Bechtel, Nicolas Clozeau, and Max Sauerbrey for their valuable suggestions and discussions. The author also acknowledges Franco Flandoli for his insights into the heuristic derivation presented in Subsection 1.2. Additionally, the author thanks Umberto Pappalettera for fruitful discussions and comments, which significantly improved the presentation of the results. Finally, the author is grateful to the anonymous referee for a careful reading of the manuscript and for providing several insightful comments.